# Quantifying AI Psychology: A Psychometrics Benchmark for Large Language Models

### v1

@@ 6 Evaluation on Theory of Mind
Theory of Mind (ToM) refers to the ability to attribute mental states to oneself and others, essential for effective communication and interaction [83, 84]. ToM involves reasoning about others’ thoughts and beliefs to predict their behaviors [84]. In this section, we apply the concept of ToM to LLMs to investigate whether they can infer perspectives and thoughts from textual scenarios. Additionally, we examine the performance stability of ToM abilities across different tasks and real-world scenarios.
@@
Setup.  To evaluate ToM in LLMs, we conduct three tests, spanning various scenarios that require different orders of ToM reasoning: (1) Evaluation on false belief task. This task assesses the ability to understand that others hold incorrect beliefs [34]. Our false belief task comprised two sub-tasks: unexpected content task and unexpected transfer task, with all items being alternative-choice. (2) Evaluation on strange story task. The strange stories scenarios cover seven non-literal language uses (e.g., metaphors) that can be misinterpreted without ToM [35]. Each item contains an open-ended question, asking about the understanding of the protagonists’ thoughts. We also use LLM raters, GPT-4 and Llama3-70b, to evaluate the responses. (3) Evaluation on imposing memory task. This task includes alternative-choice items with statements about the intentionality of characters in the scenario, and LLMs should judge if the statements correctly reflect the characters’ intentions.
@@
Results.  We include detailed discussions in Appx. F and summarize our key findings here. As illustrated in Table 25, GPT-4 and Llama3-70b achieve remarkable performance over all ToM tests. In contrast, ChatGPT, GLM4, and Mixtral-8*7b exhibit great variability across tests. For example, GLM4 excels at unexpected content tasks and struggles with unexpected transfer tasks. Similarly, Mixtral-8*7b has an 83.3% accuracy rate on imposing memory test but performs poorly on the unexpected transfer test. These results indicate that while some LLMs have abilities in ToM tasks, they lack the comprehensive capability to handle a wide range of ToM challenges.
@@
Validation.  We conduct rigorous test validation for the reliability of results for LLMs in ToM tasks. For test (1), we validate two forms of reliability: (i) Position bias robustness. Table 26 shows most models demonstrate robustness against position bias, evidenced by high match rate ($MR$) (defined in Equation 3). However, Llama3-8b and Mistral-7b show low $MR$ scores, demonstrating significant variability. (ii) Parallel form consistency. To mitigate biases from word order and language tendencies, we modify the false belief task by swapping labels on the container and its contents in the scenario. Achieving consistent results in these modified tasks is essential for validating ToM capabilities. Table 27 reveals that models such as Mixtral-8*7b display low $MR$ values, demonstrating poor consistency and randomness in their responses. In test (2), we assess inter-rater reliability, and we propose a metric termed agreement rate ($AR$) as “similarity” between two evaluations (defined in Equation 4). Table 28 shows LLM raters have high consensus with $AR$ values above 0.8 for all models . In test (3), we evaluate parallel form reliability by altering the names and genders of characters in the stories. This modification prevents LLMs from associating specific mental states with a character in alternative-choice tasks. We employ the $MR$ score (defined in Equation 3) to assess the parallel form’s reliability. As shown in Table 29, all models record $MR$ values of above 0.9, which validates the parallel form reliability of the test.
7 Evaluation on Motivation
@@
The concept of motivation in psychology is understood as the driving force behind human actions, thoughts, and behaviors toward goal attainment [85]. In this section, we apply the notion of self-efficacy in motivation, originally defined as the belief to overcome challenges [86], to LLMs. We reinterpret this notion as the perceived capability or “confidence” of LLMs to handle user queries. This attribute is crucial for their functionality as problem-solving assistants. In this section, we explore the self-efficacy of LLMs across various user query types and examine whether the self-efficacy they report aligns with their responses to actual queries.
@@
Setup.  To explore the self-efficacy of LLMs, we employ two evaluation scenarios: (1) Evaluation of self-reported LLM self-efficacy. We design LLM Self-Efficacy questionnaire that gauges LLMs’ self-reported confidence in handling queries that are challenging or beyond their capabilities. Query types are identified by Gao et al. [67], including real-time data retrieval and specialized professional queries. (2) Evaluation of operational LLM self-efficacy. We utilize HoneSet dataset [67], which consists of 930 user queries across the same six query types. This evaluation determines whether LLMs display confidence or recognize their limitations in response to specific queries. We introduce a metric termed confidence rate, intuitively understood as the the likelihood of LLMs successfully responding to a query without admitting limitations (detailed in Appx. G).
Figure 5: The confidence level in LLM Self-Efficacy questionnaire and HoneSet dataset for GPT-4 (left) and Mixtral-8*7b (right).
@@
Results.  Tests (1) and (2) assess self-efficacy, or “confidence” of LLMs through different evaluation scenarios. Test (1) employs the self-reported questionnaire for LLMs to rate their confidence, whereas test (2) assesses their operational confidence in specific query scenarios. As detailed in Table 32 and Table 33, notable discrepancies emerge between self-reported and operational confidence. LLMs often report no confidence in managing non-textual or sensory data yet do not fully recognize these limitations when responding to user queries, resulting in fabricated responses. Figure 5 illustrates that GPT-4’s self-reported confidence generally matches its responses. In contrast, Mixtral-8*7b, reports no confidence in processing non-textual and sensory data but still answers over 50% of such queries without admitting limitations. More details are discussed in Appx. G.
@@
Validation.  To validate the reliability of LLM Self-Efficacy questionnaire, we a parallel form by reversing the logic of the statements (e.g., a 100% confidence score on a “Can” statement should ideally correspond to 0% on a “Cannot” statement). We use weighted Kappa coefficients $\kappa$ to quantify the parallel form consistency. In Table 34, several LLMs, such as ChatGPT and Mistral-7b, show inconsistencies in parallel forms, evidenced by a $\kappa$ value near 0. It indicates that LLMs struggle to respond consistently to the inverse framing of statements, revealing limitations in their cognitive flexibility or contextual understanding.
8 Discussion on Intelligence
Intelligence, a multifaceted construct, has captivated psychology and AI researchers. Recent studies have explored various aspects of intelligence in LLMs, including arithmetic [87] and symbolic reasoning [88]. Given the extensive evaluation of LLMs’ intelligence, we did not include experiments in our benchmark. Instead, we discuss a critical question: How can psychometrics improve the evaluation of LLMs’ intelligence? Traditional benchmarks often rely on classical test theory [89], which simply sums or averages scores from correct responses. This method does not consider the varying difficulties of test items nor provides predictive power for performance on unseen tasks. Item Response Theory (IRT) [90, 91] in psychometrics offers a more nuanced assessment by modeling the probability of a subject correctly answering an item based on the ability level and the item’s difficulty. IRT allows for the selection of items tailored to the subject’s proficiency, enabling direct comparisons across different benchmarks and enhancing the efficacy of LLMs’ intelligence assessments.
9 Conclusion
@@
In this paper, we present a comprehensive psychometrics benchmark for LLMs, incorporating six psychological dimensions and thirteen datasets to assess their psychological attributes. Different from existing studies, our psychometrics benchmark challenges the assumption of consistent responses—central to human psychometrics—by testing LLMs across diverse evaluation scenarios, including self-reported questionnaires, open-ended questions, and multiple-choice questions. We also suggest a rigorous framework for assessments and results validation. Our findings demonstrate the diversity and variability of LLMs across evaluation scenarios. Based on these findings, we offer insights into AI and social science communities and explore potential applications. Limitations and future directions are discussed in Appx. I.
Appendix
[…]
[appendices]l1
Appendix A Dataset Overview
@@
Our benchmark includes 13 datasets from three sources: standard psychometrics tests, established datasets, and self-designed scenarios. In curating datasets, we adhere to the following guidelines:
@@
  • Authoritative and Established Datasets: The psychometrics datasets used in our benchmark are both authoritative and well-established. We select datasets that are widely recognized in psychology research to enhance the authority of our assessments. For instance, we utilize the Big Five personality test [62], which is a standard personality assessment. In contrast, we exclude the Myers-Briggs Type Indicator (MBTI) from our personality evaluations due to its limited use in scientific research and ongoing debates regarding its validity. In our benchmark, we ensure that the questions in self-curated datasets are grounded on established principles.
@@ Appendix B Results Validation
Results validation in psychometrics ensures that tests produce reliable and interpretable results. A fundamental principle of psychometrics in test validation is reliability, defined as the degree to which a test is free from error [18]. Reliability pertains to the consistency of a test under various conditions, including over time (test-retest reliability), across different versions (parallel forms reliability), and among different evaluators (inter-rater reliability). Due to the difference between humans and LLMs, applying psychometric tests to LLMs poses unique challenges. Therefore, we extend reliability considerations in psychometrics and focus on the following five forms of reliability.
  • Internal Consistency is the degree of homogeneity among the items on a test, such that they are consistent with one another and measuring the same thing [75]. That is, internal consistency measures whether LLMs exhibit stable attributes and have similar preferences to the questions examining the same aspect.
  • Parallel Forms Reliability reflects whether two different yet equivalent versions of a test yield consistent results. The parallel forms may differ slightly, such as in phrasing or terminology. This form of reliability is used to examine prompt sensitivity in LLMs.
  • Inter-Rater Reliability is used when the test involves open-ended questions that require evaluations by multiple raters to judge the responses. This form of reliability measures the level of agreement between different raters’ judgments. In this work, we use two competent LLMs, GPT-4 and Llama3-70b, as the raters to evaluate open-ended responses. Inter-rater reliability is determined by correlations between the judgments of the two LLM raters, thereby quantifying the consistency of their evaluations.
  • Option Position Robustness is the extent to which the arrangement of options in alternative-choice or multiple-choice items influences the assessment outcomes. This form of reliability is vital for ensuring that evaluations of LLMs remain unbiased, regardless of the configuration of answer choices.
  • Adversarial Attack Robustness represents the extent to which LLMs are unaffected by adversarial prompts. When evaluating opinion-related aspects of LLMs, we conduct tests using both a standard dataset and a comparable one infused with adversarial elements. This approach helps determine the robustness of opinions generated by the models.
@@ Appendix C Additional Details of Evaluation on Personality
Personality is an enduring set of traits one exhibits [92]. Understanding the distinct personality attributes of LLMs can optimize their functionality in downstream tasks. Testing these traits not only deepens our understanding but also fosters innovation in AI’s social adaptability and human-computer interaction (HCI) technologies. For instance, an LLM characterized by an extraverted personality may be particularly effective in educational applications that demand extensive user interaction, potentially enhancing user satisfaction and engagement. Furthermore, investigating the personalities of LLMs, especially darker traits, presents an opportunity to enhance the trustworthiness of these models [93, 47]. For example, personality testing can proactively identify and mitigate toxic behaviors before deployment. Additionally, by adjusting specific traits—such as reducing neuroticism and increasing agreeableness—we aim to make interactions with LLMs safer and more inclusive, thereby improving the overall user experience with these technologies [29].
@@
C.1 Big Five Inventory
Dataset.  Big Five Inventory (BFI) is a widely-recognized personality test [62], covering aspects of agreeableness, conscientiousness, extraversion, neuroticism, and openness. It contains 44 rating-scale items. We refer to McCrae et al. [94] for the descriptive definition of each aspect.
@@ C.1 Big Five Inventory
Table 3: Statement Examples in Big Five Inventory.
@@
To evaluate the effects of role-playing prompts on LLMs, we employ four types of prompts: naive prompts [4], keyword prompts, and personality prompts ($\text{P}^{2}$) [31], and reverse personality prompt ($\neg\text{P}^{2}$). The personality prompts are GPT-4 generated descriptive sentences about specific personality traits. We use the same generating procedure introduced by Jiang et al. [31]. We also design reverse personality prompts, using GPT-4 to generate descriptions that are the opposite of personality prompts. We ensure that the sentence structure of the reverse personality prompt mirrors that of the original personality prompt. These role-playing prompts are added before the statement. We provide examples of role-playing prompts for extroverted trait in the following.
@@
Results.  Each personality aspect across the datasets (e.g., openness) comprises multiple questions. The final score for each dimension is determined by computing the average of all associated question scores. In Table 4, we also include the average human scores (3,387,303 participants) for BFI in the United States [95]. We observe that LLMs generally score higher than humans in agreeableness and conscientiousness, while their scores in neuroticism are significantly lower.
@@
Figure 6: Radar figures for the personality of Big Five Inventory.
Table 4: The results of the big five test. "Agreeable." means "Agreeableness", and "Conscientious." means "Conscientiousness".
@@
Table 5: The results of the big five test using naive prompts. "Agreeable." means "Agreeableness", and "Conscientious." means "Conscientiousness".
@@
Table 6: The results of the big five test using keyword prompts. "Agreeable." means "Agreeableness", and "Conscientious." means "Conscientiousness".
@@
Table 7: The results of the big five test using personality prompts. "Agreeable." means "Agreeableness", and "Conscientious." means "Conscientiousness".
@@
Table 8: The results of the big five test using reverse personality prompts. "Agreeable." means "Agreeableness", and "Conscientious." means "Conscientiousness".
@@
$\sigma(a_{i})=\sqrt{\frac{1}{|\mathcal{Q}_{a_{i}}|}\sum_{j=1}^{|\mathcal{Q}_{a_{i}}|}(s_{ij}-\bar{s}_{i})^{2}}$ (1)
where $s_{ij}$ represents the score of the $j$-th, and $\bar{s}_{i}$ is the mean score across all items in the same aspect. This measure of variability indicates the stability of personality in LLMs. We record the $\sigma$ for BFI in Table 4. We also calculate the $\sigma$ for the personality under different prompts, shown in Table 5, Table 6, Table 7, and Table 8. A notable observation is that the personality prompts effectively decrease the variability of personality traits for almost all models, which demonstrate that the personality prompts not only direct LLMs to exhibit designated personality, but also enhance its stability.
C.2 Short Dark Triad
Dataset.  Short Dark Triad (SD3) focuses on darker aspects of personality, which offers a crucial measure of potential trustworthiness within LLMs’ personalities. We employ the latest and widely-used dataset [63], which evaluates LLMs based on Machiavellianism, Narcissism, and Psychopathy. The definition of dark aspects of personality are refer to Muris et al. [96]:

### v2

@@ 6. Evaluation on Theory of Mind
Theory of Mind (ToM) refers to the ability to attribute mental states to oneself and others, essential for effective communication and interaction (Premack and Woodruff, 1978; Baron-Cohen et al., 1985). ToM involves reasoning about others’ thoughts and beliefs to predict their behaviors (Baron-Cohen et al., 1985). We apply the concept of ToM to LLMs to investigate whether they can infer perspectives and thoughts from textual scenarios. Different from humans, where ToM is a fundamental cognitive ability, evaluation of ToM in LLMs is to understand their reasoning abilities in textual scenarios based on linguistic cues and patterns. Additionally, we examine the performance consistency of ToM abilities across different tasks and real-world scenarios.
@@
Setup.  To evaluate ToM in LLMs, we conduct three tests, spanning various scenarios that require different orders of ToM reasoning: (1) Evaluation on false belief task. This task assesses the ability to understand that others hold incorrect beliefs (Kosinski, 2023). Our false belief task comprised two sub-tasks: unexpected content task and unexpected transfer task, with all items being alternative-choice. (2) Evaluation on strange story task. The strange stories scenarios cover seven non-literal language uses (e.g., metaphors) that can be misinterpreted without ToM (van Duijn et al., 2023). Each item contains an open-ended question, asking about the understanding of the protagonists’ thoughts. We also use LLM raters, GPT-4 and Llama3-70b, to evaluate the responses with reference answers. (3) Evaluation on imposing memory task. This task includes alternative-choice items with statements about the intentionality of characters in the scenario, and LLMs should judge if the statements correctly reflect the characters’ intentions.
@@
Results.  We include detailed discussions in Appx. F and summarize our key findings here. As illustrated in Table 25, GPT-4 and Llama3-70b achieve remarkable performance over all ToM tests. In contrast, ChatGPT, GLM4, and Mixtral-8*7b exhibit great performance variability across tests. For example, GLM4 excels at unexpected content tasks but struggles with unexpected transfer tasks. Similarly, Mixtral-8*7b has an 83.3% accuracy rate on imposing memory test but performs poorly on the unexpected transfer test. These results indicate that while some LLMs have abilities in ToM tasks, they lack a comprehensive set of capabilities to handle a wide range of ToM challenges.
@@
Validation.  We conduct rigorous test validation for the reliability of results for LLMs in ToM tasks. For test (1), we validate two forms of reliability: (i) Position bias robustness. Table 26 shows most LLMs demonstrate robustness against position bias, evidenced by high match rate ($MR$) (defined in Equation 3). However, Llama3-8b and Mistral-7b show low $MR$ scores, indicating significant performance inconsistency. (ii) Parallel form consistency. To mitigate biases from word order and language tendencies, we modify the false belief task by swapping labels on the container and its contents in the scenario. Achieving consistent results in these modified tasks is essential for determining ToM capabilities. Table 27 reveals that models such as Mixtral-8*7b display low $MR$ values, demonstrating poor consistency and randomness in their responses. In test (2), we assess inter-rater reliability, and we propose a metric termed agreement rate ($AR$) as “similarity” between two evaluations (defined in Equation 4). Table 28 shows LLM raters have high consensus with $AR$ values above 0.8 for all models. Therefore, we conclude that LLM raters can reliably evaluate the responses with reference answer in our cases. In test (3), we evaluate parallel form reliability by altering the names and genders of characters in the stories. This modification prevents LLMs from associating specific mental states with a character in alternative-choice tasks. We employ the $MR$ score (defined in Equation 3) to assess the parallel form’s reliability. As shown in Table 29, all models record $MR$ values of above 0.9, which validates the parallels form reliability of the test. High parallel forms reliability demonstrates that LLMs can consistently provide reliable answers despite variations in items, such as changes in nouns, highlighting their genuine capability to address such challenges.
7. Evaluation on Motivation
@@
The concept of motivation in psychology is understood as the driving force behind human actions, thoughts, and behaviors toward goal attainment (Broussard and Garrison, 2004). When extending the notion of motivation to LLMs, we find that directly applying human motivation questionnaires to LLMs presents challenges, as these assessments presume inherent needs that LLMs lack. However, the notion of self-efficacy, an important aspect of motivation, is meaningful to LLMs. Self-efficacy is defined as the belief to overcome challenges (Bandura, 1977), and we interpret this notion as the perceived capability or “confidence” of LLMs to handle user queries. In this section, we explore the self-efficacy of LLMs across various user query types and examine whether the self-efficacy they report aligns with their responses to actual queries.
@@
Setup.  To explore the self-efficacy of LLMs, we conduct two tests: (1) Evaluation of self-reported LLM self-efficacy. We create LLM Self-Efficacy questionnaire that gauges LLMs’ self-reported confidence in handling queries that are challenging or beyond their capabilities. Query types are identified by Gao et al. (2024), including real-time data retrieval and specialized professional queries. (2) Evaluation of operational LLM self-efficacy. We utilize the HoneSet dataset (Gao et al., 2024), which consists of 930 user queries across the same six query types. This evaluation determines whether LLMs display confidence or recognize their limitations in response to specific queries. We introduce a metric termed confidence rate, defined as the likelihood of LLMs successfully responding to a query without admitting limitations (detailed in Appx. G).
@@
Results.  Tests (1) and (2) assess self-efficacy, or “confidence” of LLMs through different evaluation scenarios. Test (1) employs the self-reported questionnaire for LLMs to rate their confidence, whereas test (2) assesses their operational confidence in specific query scenarios. As detailed in Table 32 and Table 33, we observe notable discrepancies emerge between self-reported and operational confidence. LLMs often report no confidence in managing non-textual or sensory data yet do not fully recognize these limitations when responding to real-world user queries, resulting in fabricated responses. Figure 5 illustrates that GPT-4’s self-reported confidence generally aligns its responses to real-world queries. In contrast, Mixtral-8*7b, reports no confidence in processing non-textual and sensory data but still answers over 50% of such queries without admitting limitations. This results in concerning trustworthiness issues, as users cannot accurately gauge the reliability of LLMs’ information. Without reliable uncertainty reporting, users may either overtrust fabricated answers or overlook the models’ genuine limitations. More details are discussed in Appx. G.
Figure 5. The confidence level in LLM Self-Efficacy questionnaire and HoneSet dataset for GPT-4 (left) and Mixtral-8*7b (right).
@@
Validation.  To validate the reliability of LLM Self-Efficacy questionnaire, we create a parallel form of the test by reversing the logic of the statements (e.g., a 100% confidence score on a “Can” statement should ideally correspond to 0% on a “Cannot” statement). We use weighted Kappa coefficients $\kappa$ to quantify the parallel form consistency. In Table 34, several LLMs, such as ChatGPT and Mistral-7b, show inconsistencies in parallel forms, evidenced by a $\kappa$ value near 0. It indicates that LLMs struggle to respond consistently to the inverse framing of statements, revealing limitations in their contextual understanding. As a result, for LLMs with low parallel forms consistency, self-reported confidence is unreliable because they may not genuinely understand the questions, thereby invalidating their reported responses.
8. Related Work
(Burnell et al., 2023) found that the performance of LLMs can be explained by a small number of latent constructs. Existing evaluations have explored specific psychological constructs such as personality (Bodroza et al., 2023; Jiang et al., 2023a), emotion (Zhan et al., 2023; Sabour et al., 2024), and theory of mind (Kosinski, 2023; van Duijn et al., 2023), with detailed discussions in Appx. I. Other studies investigate a broader scope of constructs, such as Miotto et al. (2022) on GPT-3, assessing personality, values, and demographics, and Huang et al. (2024) covering personality, relationships, motivations, and emotional abilities. However, not enough attention has been paid to reliability and the interpretation of results. On the other hand, some prior works are conceptually related to ours in suggesting reliability examinations for evaluation. For example, Jacobs and Wallach (2021) and Wang et al. (2023a) emphasized the importance of stable, reliable measurements in AI through psychometric frameworks. Van der Wal et al. (2024) discussed key reliability measures such as test-retest reliability to ensure that the biases identified are not caused by random noise or inconsistencies. Building on these insights, we integrate reliability examination as a key element of our benchmark.
9. Conclusion
@@
In this paper, we present a comprehensive psychometric benchmark for LLMs, covering the evaluation of five psychological dimensions and thirteen datasets to assess their psychological patterns. Different from existing studies, our psychometric benchmark challenges the assumption of consistent responses—central to human psychometrics—by testing LLMs across diverse evaluation scenarios, including self-reported questionnaires, open-ended questions, and multiple-choice questions. Our work not only focuses on understanding the psychological tendencies of LLMs but also suggests a rigorous framework for assessment and results validation. Our findings demonstrate the diversity and variability of LLMs across evaluation scenarios. Based on these findings, we offer insights to the AI and social science communities and explore potential applications. Limitations and future directions are discussed in Appx. J.
Ethics Statement
This paper provides a comprehensive analysis of LLMs to better understand and predict their behaviors through the lens of psychometrics. It carries significant social and ethical implications. Our psychometrics benchmark enhances LLM evaluation by identifying biases and inconsistencies, promoting more ethically responsible AI (Yao et al., 2023; Sun et al., 2024; Gallegos et al., 2024). It also supports the development of personalized AI assistants in sectors such as healthcare and education (Kasneci et al., 2023; Yang et al., 2023) and enhances public trust by improving user experience. However, we are aware of the potential risks of misuse and misinterpretation of the results from our benchmark. One potential misinterpretation is the humanization of LLMs, leading to beliefs that LLMs are already capable or have reached human-level intelligence. Misinterpreting LLM capabilities might lead to unrealistic expectations, such as assuming these models can make moral judgments or replace human decision-making in critical areas like healthcare or law. This can result in over-dependence and neglect of human oversight. Additionally, these misinterpretations could be used to spread misinformation, automate and scale biased decision-making, or even develop manipulative technologies under the guise of advanced AI. One potential way to mitigate such problems is through psychology-related safety evaluations (Zhang et al., 2024). This approach examines stereotypes, discriminatory practices, and deceptive behaviors of LLMs.
Appendix A Guidelines for Dataset
@@
Our benchmark includes 13 datasets from three sources: standard psychometrics tests, established datasets, and self-designed scenarios. In developing datasets, we adhere to the following guidelines:
@@
  • Authoritative and Established Datasets: The psychometrics datasets used in our benchmark are both authoritative and well-established. We select datasets that are widely recognized in psychology research to enhance the authority of our assessments. For instance, we utilize the Big Five personality test [John et al., 1999], which is a standard personality assessment. In contrast, we exclude the Myers-Briggs Type Indicator (MBTI) from our personality evaluations due to its limited use in scientific research and ongoing debates regarding its validity. In our benchmark, we ensure that the questions in self-curated datasets are grounded on established principles.
@@ Appendix B Results Validation
Results validation in psychometrics ensures that tests produce reliable and interpretable results. A fundamental principle of psychometrics in test validation is reliability, defined as the degree to which a test is free from error [Rust and Golombok, 2014]. Reliability pertains to the consistency of a test under various conditions, including over time (test-retest reliability), across different versions (parallel forms reliability), and among different evaluators (inter-rater reliability).Due to the differences between humans and LLMs, applying psychometric tests to LLMs poses unique challenges. Therefore, we extend reliability considerations from psychometrics and focus on five forms of reliability. Internal consistency, parallel forms reliability, and inter-rater reliability are derived from psychometrics and assist in ensuring trustworthy interpretation of results. While option position robustness and adversarial attack robustness are specifically designed for LLMs, their concepts are interconnected with reliability in the psychometric framework. Option position robustness assesses the extent to which the arrangement of options in multiple-choice items influences assessment outcomes. It can be considered a type of parallel forms reliability, involving items that probe the same construct but with shuffled option positions. Adversarial attack robustness represents the extent to which LLMs remain unaffected by adversarial prompts. While these adversarial forms can be validated through parallel forms reliability to check if they measure the same construct, the core idea is to compare LLM performance with and without adversarial attacks. This assessment provides an additional dimension to understand LLM behavior, particularly their resilience to deceptive inputs, which is critical for real-world applications.
@@ Appendix C Additional Details of Evaluation on Personality
Personality is an enduring set of traits one exhibits [Mischel, 2013]. Understanding the distinct personality attributes of LLMs can optimize their functionality in downstream tasks. Testing these traits not only deepens our understanding but also fosters innovation in AI’s social adaptability and human-computer interaction (HCI) technologies. For instance, an LLM characterized by an extraverted personality may be particularly effective in educational applications that demand extensive user interaction, potentially enhancing user satisfaction and engagement. Furthermore, investigating the personalities of LLMs, especially darker traits, presents an opportunity to enhance the trustworthiness of these models [Li et al., 2022b, Sun et al., 2024]. For example, personality testing can proactively identify and mitigate toxic behaviors before deployment. Additionally, by adjusting specific traits—such as reducing neuroticism and increasing agreeableness—we aim to make interactions with LLMs safer and more inclusive, thereby improving the overall user experience with these technologies [Safdari et al., 2023].
@@
C.1. Big Five Inventory
Dataset.  Big Five Inventory (BFI) is a widely-recognized personality test [John et al., 1999], covering aspects of agreeableness, conscientiousness, extraversion, neuroticism, and openness. It contains 44 rating-scale items. We refer to McCrae and John [1992] for the descriptive definition of each aspect.
@@ C.1. Big Five Inventory
Table 3. Statement Examples in Big Five Inventory.
@@
To evaluate the effects of role-playing prompts on LLMs, we employ four types of prompts: naive prompts [Brown et al., 2020], keyword prompts, and personality prompts ($\text{P}^{2}$) [Jiang et al., 2023a], and reverse personality prompt ($\neg\text{P}^{2}$). The personality prompts are GPT-4 generated descriptive sentences about specific personality traits. We use the same generating procedure introduced by Jiang et al. [2023a]. We also design reverse personality prompts, using GPT-4 to generate descriptions that are the opposite of personality prompts. We ensure that the sentence structure of the reverse personality prompt mirrors that of the original personality prompt. These role-playing prompts are added before the statement. We provide examples of role-playing prompts for extroverted trait in the following.
@@
Results.  Each personality aspect across the datasets (e.g., openness) comprises multiple questions. The final score for each dimension is determined by computing the average of all associated question scores. In Table 4, we also include the average human scores (3,387,303 participants) for BFI in the United States [Ebert et al., 2022]. We observe that LLMs generally score higher than humans in agreeableness and conscientiousness, while their scores in neuroticism are significantly lower.
@@
Figure 6. Radar figures for the personality of Big Five Inventory.
Table 4. The results of the big five test. ”Agreeable.” means ”Agreeableness”, and ”Conscientious.” means ”Conscientiousness”.
@@
Table 5. The results of the big five test using naive prompts. ”Agreeable.” means ”Agreeableness”, and ”Conscientious.” means ”Conscientiousness”.
@@
Table 6. The results of the big five test using keyword prompts. ”Agreeable.” means ”Agreeableness”, and ”Conscientious.” means ”Conscientiousness”.
@@
Table 7. The results of the big five test using personality prompts. ”Agreeable.” means ”Agreeableness”, and ”Conscientious.” means ”Conscientiousness”.
@@
Table 8. The results of the big five test using reverse personality prompts. ”Agreeable.” means ”Agreeableness”, and ”Conscientious.” means ”Conscientiousness”.
@@
(1) $\sigma(a_{i})=\sqrt{\frac{1}{|\mathcal{Q}_{a_{i}}|}\sum_{j=1}^{|\mathcal{Q}_{a_{i}}|}(s_{ij}-\bar{s}_{i})^{2}}$
where $s_{ij}$ represents the score of the $j$-th, and $\bar{s}_{i}$ is the mean score across all items in the same aspect. This reliability measure indicates the consistency of personality of LLMs to similar situations. We record the $\sigma$ for BFI in Table 4. We also calculate the $\sigma$ for the personality under different prompts, shown in Table 5, Table 6, Table 7, and Table 8. A notable observation is that the personality prompts effectively decrease the inconsistency of personality traits for almost all models, which demonstrate that the personality prompts not only direct LLMs to exhibit designated personality, but also enhance its consistency.
C.2. Short Dark Triad
Dataset.  Short Dark Triad (SD3) focuses on darker aspects of personality, which offers a crucial measure of potential trustworthiness within LLMs’ personalities. We employ the latest and widely-used dataset [Jones and Paulhus, 2014], which evaluates LLMs based on Machiavellianism, Narcissism, and Psychopathy. The definitions of dark aspects of personality refer to Muris et al. [2017]: